# A tuning-free and scalable method for joint graphical model estimation with sharper bounds

### v2

@@ 2.2 The MIGHT method
This subsection introduces a node-wise Multi-task Iterative Graphical Hard Thresholding (MIGHT) to solve the scaled model (4). The key to this method is a two-step thresholding operator $\mathcal{T}_{\lambda,s_{0}}:\mathbb{R}^{(p-1)K}\to\mathbb{R}^{(p-1)K}$ inspired by Zhang et al. (2024), which can be constructed from the following definitions.
@@
    For a vector $\bm{\beta}=\left(\left(\bm{\beta}^{(1)}\right)^{\top},\cdots,\left(\bm{\beta}^{(K)}\right)^{\top}\right)^{\top}\in\mathbb{R}^{(p-1)K}$, where each subvector $\bm{\beta}^{(k)}\in\mathbb{R}^{p-1}$ corresponds to the $k$-th dataset, we define the edge-wise thresholding operator $\mathcal{T}_{\lambda}^{(1)}:\mathbb{R}^{(p-1)K}\to\mathbb{R}^{(p-1)K}$ as:
    $\left\{\mathcal{T}_{\lambda}^{(1)}(\bm{\beta})\right\}_{i}^{(k)}=\bm{\beta}_{i}^{(k)}\cdot\mathbf{1}\left(|\bm{\beta}_{i}^{(k)}|\geq\lambda\right)\in\mathbb{R},\text{ for each }(i,k)\in[p-1]\times[K],$
    where $\bm{\beta}_{i}^{(k)}$ represents the $i$-th entry of $\bm{\beta}^{(k)}$.
@@
    We define the neighbor-wise thresholding operator $\mathcal{T}_{\lambda,s_{0}}^{(2)}:\mathbb{R}^{(p-1)K}\to\mathbb{R}^{(p-1)K}$ as:
    $\left\{\mathcal{T}_{\lambda,s_{0}}^{(2)}(\bm{\beta})\right\}^{(k)}_{i}=\bm{\beta}^{(k)}_{i}\cdot\mathbf{1}\left\{\sum_{k^{\prime}\in[K]}\left(\bm{\beta}^{(k^{\prime})}_{i}\right)^{2}\geq s_{0}\lambda^{2}\right\}\in\mathbb{R},\text{for each }(i,k)\in[p-1]\times[K].$
Then, the two-step hard thresholding operator can be described as $\mathcal{T}_{\lambda,s_{0}}:=\mathcal{T}_{\lambda,s_{0}}^{(2)}\circ\mathcal{T}_{\lambda}^{(1)}$, i.e., $\mathcal{T}_{\lambda,s_{0}}(\bm{\beta}):=\mathcal{T}_{\lambda,s_{0}}^{(2)}\left(\mathcal{T}_{\lambda}^{(1)}(\bm{\beta})\right)$. Specifically, in the multi-task learning problem (4) with respect to node $j$, for each potential neighbor $i\in[p]\setminus\{j\}$, if $\sum_{k\in[K]}\left\{\left(\hat{\bm{\beta}}_{i}^{(k)}\right)^{2}\cdot\mathbf{1}\left(|\hat{\bm{\beta}}_{i}^{(k)}|\geq\lambda\right)\right\}\geq s_{0}\lambda^{2}$ holds, node $i$ is selected as a neighbor of node $j$, and its edge takes the thresholded value
@@
This thresholding mechanism simultaneously controls the edge-sparsity and neighbor-similarity across multiple graphical models.
We propose Algorithm 1 to solve the scaled model (4). This algorithm is based on the least squares loss $\mathcal{L}(\bm{\beta}_{\setminus j})=\frac{2}{N}\left\|X_{j}-Z_{\setminus j}\bm{\beta}_{\setminus j}\right\|_{2}^{2}$ and iteratively updates parameters by applying the operator $\mathcal{T}_{\lambda,s_{0}}$. The algorithm consists of two stages: the dynamic thresholding iteration and the fixed thresholding iteration. In the dynamic thresholding stage, the thresholding parameter $\lambda_{t}$ decreases geometrically from an initial value $\lambda_{0}$ to a pre-specified fixed value $\lambda_{\infty}$, which effectively balances Type I and Type II errors and yields an initial estimation of $\bm{\beta}_{\setminus j}$. In the fixed thresholding stage, we construct an appropriate thresholding parameter $\lambda_{\text{fix}}$, which remains fixed in subsequent iterations. This stage refines the initial estimation, ensuring selection consistency and asymptotic properties.
Algorithm 1
0:  $Z_{\setminus j},~{}X_{j},~{}\kappa,~{}s_{0}$
1:  Initialize $t=0,~{}\lambda_{\infty}=C_{1}\sqrt{\frac{\log K+(1/s_{0})\log p}{N}}$, $\hat{\bm{\beta}}_{\setminus j}^{0}=\mathbf{0}_{(p-1)K}$ and $N=\sum_{k\in[K]}n_{k}$
2:  Initialize $\lambda_{0}=C_{2}\left(\sqrt{\frac{\log(pK)}{N}}+\|\frac{1}{N}Z_{\setminus j}^{\top}X_{j}\|_{\infty}\right)$
@@
4:     $\hat{\bm{\beta}}_{\setminus j}^{t+1}=\mathcal{T}_{\lambda_{t},s_{0}}\left(\hat{\bm{\beta}}_{\setminus j}^{t}+\frac{1}{N}Z_{\setminus j}^{\top}(X_{j}-Z_{\setminus j}\hat{\bm{\beta}}_{\setminus j}^{t})\right)$ // Dynamic thresholding stage
@@
8:  $\hat{s}=\max\left\{1,~{}\left|\bigcup_{k\in[K]}\text{supp}\left(\left(\hat{\bm{\beta}}_{\setminus j}^{t}\right)^{(k)}\right)\right|\right\}$ and $\lambda_{\text{fix}}=C_{3}\sqrt{\frac{\log(\hat{s}K)+(1/s_{0})\log p}{N}}$
9:  while $t\leq\left\lceil\frac{\log(\lambda_{0}/\lambda_{\infty})}{\log(1/\kappa)}\right\rceil+C_{4}\log N,~{}$ do
10:     $\hat{\bm{\beta}}_{\setminus j}^{t+1}=\mathcal{T}_{\lambda_{\text{fix}},s_{0}}\left(\hat{\bm{\beta}}_{\setminus j}^{t}+\frac{1}{N}Z_{\setminus j}^{\top}(X_{j}-Z_{\setminus j}\hat{\bm{\beta}}_{\setminus j}^{t})\right)$ // Fixed thresholding stage
11:     $t=t+1$
12:  end while
12:  $\hat{\bm{\beta}}_{\setminus j}^{t}$
Building on Algorithm 1, we outline the node-wise Multi-task Iterative Graphical Hard Thresholding (MIGHT) Algorithm 2 for the joint graphical estimation. For each node $j$, Algorithm 2 firstly constructs a column-scaled multi-task model (4), then applies Algorithm 1 to obtain a node-wise estimator, and finally recovers the joint model through the relationship (3). Notably, Algorithm 2 allows parallel computation across covariates (Lines $6-13$), ensuring high computational efficiency.
@@
1:  for $k=1$ to $K,~{}$ do
2:     $\hat{\Sigma}^{(k)}=\frac{1}{n_{k}}\left(X^{(k)}\right)^{\top}X^{(k)}$ and $\hat{\Gamma}^{(k)}=\text{diag}\left(\{\hat{\Sigma}^{(k)}_{ii}\}_{i\in[p]}\right)$
3:     $Z^{(k)}=\sqrt{\frac{C_{0}N}{n_{k}}}X^{(k)}\left(\hat{\Gamma}^{(k)}\right)^{-1/2}$ // The column-wise scaling
4:  end for
5:  for $j=1$ to $p,~{}$ do
6:     $X_{j}=\left(\left(X_{j}^{(k)}\right)^{\top},\cdots,\left(X_{j}^{(k)}\right)^{\top}\right)^{\top}$ and $Z_{\setminus j}=\text{diag}\left(\{Z^{(k)}_{\setminus j}\}_{k\in[K]}\right)$
7:     $\hat{\bm{\beta}}_{\setminus j}=\mathtt{Algorithm1}(Z_{\setminus j},X_{j},\kappa,\hat{s}_{0,j})$, with $\hat{s}_{0,j}$ tuned by (5) // Multi-task linear model
8:     for $k=1$ to $K,~{}$ do
@@
10:        $\hat{\Theta}_{\setminus j,j}^{(k)}=-\sqrt{\frac{C_{0}N}{n_{k}}}\hat{\Theta}_{j,j}^{(k)}\left(\hat{\Gamma}_{\setminus j,\setminus j}^{(k)}\right)^{-1/2}\hat{\bm{\beta}}_{\setminus j}^{(k)}$ // Rescaling
@@ 2.3 Miscellaneous setups
To ensure the adaptability, we employ a variant of the Birgé-Massart criterion, as introduced by Zhang et al. (2024) to determine the neighbor-similarity degree $s_{0}$ in Algorithm 1. Concretely, for each $j\in[p]$, let $\hat{\bm{\beta}}_{\setminus j}(s_{0})$ denote the estimation obtained under parameter $s_{0}$ in Algorithm 1. We define $\hat{s}(s_{0}):=\max\left\{1,~{}\left|\bigcup_{k\in[K]}\text{supp}\left(\hat{\bm{\beta}}_{\setminus j}^{(k)}(s_{0})\right)\right|\right\}$ as the estimated number of covariates connected to node $j$, and $\hat{A}(s_{0}):=\left\|\hat{\bm{\beta}}_{\setminus j}(s_{0})\right\|_{0}$ as the total number of selected edges among all $K$ graphs. We select $\hat{s}_{0,j}$ by the following information criterion:
$\hat{s}_{0,j}=\arg\min_{s_{0}}\left\{\log\left(\frac{1}{N}\left\|X_{j}-Z_{\setminus j}\hat{\bm{\beta}}_{\setminus j}(s_{0})\right\|_{2}^{2}\right)+\frac{C}{N}\left[\hat{s}(s_{0})\log p+\hat{A}(s_{0})\log\left(K\cdot\hat{s}(s_{0})\right)\right]\right\}.$ (5)
Considering the symmetry of the precision matrix, we adopt the minimum symmetrization approach (Cai et al., 2011, 2016; Shu et al., 2024):
@@ 3 Theoretical analysis
Assume the joint graphical model $\left(\Theta^{(1)},\cdots,\Theta^{(K)}\right)$ has the joint $(s_{0},s)$ structure, i.e., has both $s_{0}$ neighbor-similarity and $s$ edge-sparsity, with $11\leq s_{0}\leq K,~{}1\leq s\leq p-$ 1.
Assumption 1 requires that the observations follow from Gaussian distributions, and the conclusions derived in this section remain valid under sub-Gaussian cases. Assumption 2 excludes the cases of singular or nearly singular precision matrices and ensures that the variance in each nodewise regression, i.e., $\text{Var}(\epsilon_{j}^{(k)})=1/\Theta_{j,j}^{(k)}$, neither diverges nor approaches zero for every $(j,k)\in[p]\times[K]$. This assumption is common in studies on graphical models (Ren et al., 2015; Ma and Michailidis, 2016). Assumption 3 characterizes the structure of the joint graphical model, as we introduced in Definition 1 and 2, which indicates $|S_{j}|\leq s$ and $\sum_{k\in[K]}|S_{j}^{(k)}|\leq ss_{0}$.
@@ 3.1 Estimation error bound
The following theorem provides column-wise bounds on the estimation error. For ease of display, for every $j\in[p]$ and $k\in[K]$, we define $\hat{S}_{j}^{(k)}:=\text{supp}\left(\hat{\Theta}_{\setminus j,j}^{(k)}\right)$ as the estimated neighbor set of node $j$ in the $k$-th graphical model.
@@
Assume Assumptions 1-3 hold and $n\gtrsim s\log p+ss_{0}\log K$. Then, by using the MIGHT Algorithm 2, we simultaneously have the following properties with probability at least $1-C_{3}e^{-C_{4}\log(pK)}$:
@@
    $\max_{j\in[p]}\sum_{k=1}^{K}\frac{n_{k}}{N}\left\|\hat{\Theta}^{(k)}_{\cdot,j}-\Theta^{(k)}_{\cdot,j}\right\|_{2}^{2}\lesssim$ $\frac{K+s\log p+ss_{0}\log(sK)}{N}.$ (6)
@@
    Sparse neighbor selection and edge selection:
    $\max_{j\in[p]}\left|\bigcup_{k\in[K]}\hat{S}_{j}^{(k)}\right|=O(s),\text{ and }\max_{j\in[p]}\sum_{k\in[K]}\left|\hat{S}_{j}^{(k)}\right|=O(ss_{0}).$
The joint estimation error in (6) consists of three components: (i) estimating the $K$ non-zero diagonal entries $\hat{\Theta}_{j,j}^{(k)}$, with rate $K/N$; (ii) estimating $s$ neighbors from $p-1$ nodes, with rate $(s\log p)/N$; and (iii) estimating $ss_{0}$ edges from $sK$ potential edges, with rate $(ss_{0}\log(sK))/N$. Additionally, from (6) we obtain the joint estimation error bound in Frobenius norm as
$\sum_{k=1}^{K}\frac{n_{k}}{N}\left\|\hat{\Theta}^{(k)}-\Theta^{(k)}\right\|_{F}\lesssim\sqrt{\frac{p\left(K+s\log p+ss_{0}\log(sK)\right)}{N}}.$ (7)
This result offers a more refined error bound than existing joint estimation methods (Cai et al., 2016; Wang et al., 2017). Furthermore, if $s_{0}=o(K/\log(sK))$, i.e., the similarity among the $K$ graphs is relatively weak, (7) indicates that our estimator achieves higher accuracy than the estimator relying on prior support structure information (Ma and Michailidis, 2016). This demonstrates the superior accuracy and adaptability of our method.

### v3

@@ 2.2 The MIGHT method
This subsection introduces a nodewise Multi-task Iterative Graphical Hard Thresholding (MIGHT) to solve the scaled model (4). The key to this method is a two-step thresholding operator $\mathcal{T}_{\lambda,s_{0}}:\mathbb{R}^{(p-1)K}\to\mathbb{R}^{(p-1)K}$ inspired by Zhang et al. (2024), which can be constructed from the following definitions.
@@
    For a vector $\bm{\beta}=\left(\left(\bm{\beta}^{(1)}\right)^{\top},\cdots,\left(\bm{\beta}^{(K)}\right)^{\top}\right)^{\top}\in\mathbb{R}^{(p-1)K}$, where each subvector $\bm{\beta}^{(k)}\in\mathbb{R}^{p-1}$ corresponds to the $k$-th dataset, we define the edge-wise thresholding operator $\mathcal{T}_{\lambda}^{(1)}:\mathbb{R}^{(p-1)K}\to\mathbb{R}^{(p-1)K}$ as: { T_λ^(1)(β)}_i^(k) = β_i^(k) ⋅1 (|β_i^(k)| ≥λ) , for each $(i,k)\in[p-1]\times[K]$, where $\bm{\beta}_{i}^{(k)}$ represents the $i$-th entry of $\bm{\beta}^{(k)}$.
@@
    We define the neighbor-wise thresholding operator $\mathcal{T}_{\lambda,s_{0}}^{(2)}:\mathbb{R}^{(p-1)K}\to\mathbb{R}^{(p-1)K}$ as: { T_λ,s_0^(2)(β)}^(k)_i = β^(k)_i ⋅1 { ∑_k’ ∈[K] ( β^(k’)_i )^2 ≥s_0 λ^2 } , for each $(i,k)\in[p-1]\times[K]$.
The two-step hard thresholding can be described as $\mathcal{T}_{\lambda,s_{0}}:=\mathcal{T}_{\lambda,s_{0}}^{(2)}\circ\mathcal{T}_{\lambda}^{(1)}$, i.e., $\mathcal{T}_{\lambda,s_{0}}(\bm{\beta}):=\mathcal{T}_{\lambda,s_{0}}^{(2)}\left(\mathcal{T}_{\lambda}^{(1)}(\bm{\beta})\right)$. Specifically, in the multi-task learning problem (4) with respect to node $j$, for each potential neighbor $i\in[p]\setminus\{j\}$, if $\sum_{k\in[K]}\left\{\left(\hat{\bm{\beta}}_{i}^{(k)}\right)^{2}\cdot\mathbf{1}\left(|\hat{\bm{\beta}}_{i}^{(k)}|\geq\lambda\right)\right\}\geq s_{0}\lambda^{2}$ holds, node $i$ is selected as a neighbor of node $j$, and its edge takes the thresholded value
@@
This thresholding mechanism simultaneously controls the edge-sparsity and neighbor-similarity across multiple graphs.
We propose Algorithm 1 to solve the scaled model (4). Starting from the least-squares loss $\mathcal{L}(\bm{\beta}_{\setminus j})=\frac{2}{N}\left\|X_{j}-Z_{\setminus j}\bm{\beta}_{\setminus j}\right\|_{2}^{2}$, in each iteration round the algorithm performs a gradient-descent update and then applies the operator $\mathcal{T}_{\lambda_{t},s_{0}}$ to enforce the desired joint structure. It consists of two stages: the dynamic thresholding iteration and the fixed thresholding iteration. In the dynamic stage (lines 3-7), the thresholding parameter $\lambda_{t}$ decreases geometrically from an initial value $\lambda_{0}$ to a pre-specified fixed value $\lambda_{\infty}$, which effectively balances Type I and Type II errors and yields an initial estimation of $\bm{\beta}_{\setminus j}$. Here, $\kappa\in(0,1)$ is a fixed decreasing rate, typically chosen as 0.9 in practice, and the explicit form of constants $C_{1}^{\text{alg}},C_{2}^{\text{alg}}$ will be given in Theorem 1. In the fixed stage (lines 10-13), we construct an appropriate thresholding parameter $\lambda_{\text{fix}}$, which remains fixed in subsequent iterations. This stage refines the initial estimation, ensuring selection consistency and asymptotic properties.
Algorithm 1 Nodewise Procedure
0:  $Z_{\setminus j},~X_{j},~\kappa,~s_{0,j}$
1:  Initialize $t=0,~\lambda_{\infty}=C_{1}^{\text{alg}}\sqrt{\frac{\log K+(1/s_{0,j})\log p}{N}}$, $\hat{\bm{\beta}}_{\setminus j}^{0}=\mathbf{0}_{(p-1)K}$ and $N=\sum_{k\in[K]}n_{k}$
2:  Initialize $\lambda_{0}=C_{2}^{\text{alg}}$
@@
4:   $\hat{\bm{\beta}}_{\setminus j}^{t+1}=\mathcal{T}_{\lambda_{t},s_{0}}\left(\hat{\bm{\beta}}_{\setminus j}^{t}+\frac{1}{N}Z_{\setminus j}^{\top}(X_{j}-Z_{\setminus j}\hat{\bm{\beta}}_{\setminus j}^{t})\right)$
@@
8:  $\hat{s}=\max\left\{1,~\left|\bigcup_{k\in[K]}\text{supp}\left(\left(\hat{\bm{\beta}}_{\setminus j}^{t}\right)^{(k)}\right)\right|\right\}$
9:  $\lambda_{\text{fix}}=C_{1}^{\text{alg}}\sqrt{\frac{\log(\hat{s}K)+(1/s_{0,j})\log p}{N}}$
10:  while $t\leq\left\lceil\frac{\log(\lambda_{0}/\lambda_{\infty})}{\log(1/\kappa)}\right\rceil+C_{4}\log N,$~ do
11:   $\hat{\bm{\beta}}_{\setminus j}^{t+1}=\mathcal{T}_{\lambda_{\text{fix}},s_{0,j}}\left(\hat{\bm{\beta}}_{\setminus j}^{t}+\frac{1}{N}Z_{\setminus j}^{\top}(X_{j}-Z_{\setminus j}\hat{\bm{\beta}}_{\setminus j}^{t})\right)$
12:   $t=t+1$
13:  end while
13:  $\hat{\bm{\beta}}_{\setminus j}^{t}$
Building on Algorithm 1, we outline the nodewise Multi-task Iterative Graphical Hard Thresholding (MIGHT) Algorithm 2 for the joint graphical estimation. For each node $j$, Algorithm 2 firstly constructs a column-scaled multi-task model (4), then applies Algorithm 1 to obtain a nodewise estimator, and finally recovers the joint model through the relationship (3). Notably, Algorithm 2 allows parallel computation across covariates (lines 6126-12), ensuring high computational efficiency.
@@
1:  for $k=1$ to K,K,~ do
2:   $\hat{\Sigma}^{(k)}=\frac{1}{n_{k}}\left(X^{(k)}\right)^{\top}X^{(k)}$
3:   $\hat{\Gamma}^{(k)}=\text{diag}\left(\{\hat{\Sigma}^{(k)}_{ii}\}_{i\in[p]}\right)$
4:   $Z^{(k)}=\sqrt{\frac{C_{0}N}{n_{k}}}X^{(k)}\left(\hat{\Gamma}^{(k)}\right)^{-1/2}$ //Column scaling
5:  end for
6:  for $j=1$ to p,p,~ do
7:   $\hat{\bm{\beta}}_{\setminus j}=\mathtt{Algorithm1}(Z_{\setminus j},X_{j},\kappa,\hat{s}_{0,j})$, with $\hat{s}_{0,j}$ tuned by (5) // Multi-task model
8:   for $k=1$ to K,K,~ do
@@
10:    $\hat{\Theta}_{\setminus j,j}^{(k)}=-\sqrt{\frac{C_{0}N}{n_{k}}}\hat{\Theta}_{j,j}^{(k)}\left(\hat{\Gamma}_{\setminus j,\setminus j}^{(k)}\right)^{-1/2}\hat{\bm{\beta}}_{\setminus j}^{(k)}$
@@ 2.3 Miscellaneous setups
To ensure the adaptability, we employ a variant of the Birgé-Massart criterion, as introduced by Zhang et al. (2024), to determine the neighbor-similarity degree $s_{0,j}$ in Algorithm 1. Concretely, for each $j\in[p]$, let $\hat{\bm{\beta}}_{\setminus j}(s_{0,j})$ denote the estimation obtained under parameter $s_{0,j}$ in Algorithm 1. We define $\hat{s}(s_{0,j}):=\max\left\{1,~\left|\bigcup_{k\in[K]}\text{supp}\left(\hat{\bm{\beta}}_{\setminus j}^{(k)}(s_{0,j})\right)\right|\right\}$ as the estimated number of neighbors to node $j$, and $\hat{A}(s_{0,j}):=\left\|\hat{\bm{\beta}}_{\setminus j}(s_{0,j})\right\|_{0}$ as the total number of selected edges (to node $j$) among all $K$ graphs. We select $\hat{s}_{0,j}$ by the following information criterion:
$\hat{s}_{0,j}=$ $\arg\min_{s_{0,j}}\left\{\log\left(\frac{1}{N}\left\|X_{j}-Z_{\setminus j}\hat{\bm{\beta}}_{\setminus j}(s_{0,j})\right\|_{2}^{2}\right)\right.+\left.\frac{C}{N}\left[\hat{s}(s_{0,j})\cdot\log p+\hat{A}(s_{0,j})\cdot\log\left(K\cdot\hat{s}(s_{0})\right)\right]\right\}.$ (5)
Considering the symmetry of the precision matrix, we adopt the minimum symmetrization approach Cai et al. (2011, 2016); Shu et al. (2024):
@@ 3 Theoretical analysis
Assume the joint graphical model $\left(\Theta^{(1)},\cdots,\Theta^{(K)}\right)$ has the joint $(s_{j},s_{0,j})_{j\in[p]}$ structure, i.e., has both $\{s_{j}\}_{j\in[p]}$ local edge-sparsity and $\{s_{0,j}\}_{j\in[p]}$ local neighbor-similarity, with $1\leq s_{0,j}\leq K,~1\leq s_{j}\leq p-1$ for each $j\in[p]$.
Assumption 1 requires that the observations follow from Gaussian distributions, and this condition can be relaxed to sub-Gaussian cases. Assumption 2 excludes singular or nearly singular precision matrices and ensures that the variance in each nodewise regression, i.e., $\text{Var}(\epsilon_{j}^{(k)})=1/\Theta_{j,j}^{(k)}$, neither diverges nor approaches zero. This assumption is common in studies on graphical models Ren et al. (2015); Ma and Michailidis (2016). Assumption 3 characterizes the structure of the joint graphs, as we introduced in Definition 2 and 1, which indicates $|S_{j}|=s_{j}$ and $\sum_{k\in[K]}|S_{j}^{(k)}|=s_{j}s_{0,j}$ for every $j\in[p]$.
@@ 3.1 Estimation error bound
The following theorem provides column-wise bounds on the estimation error. For ease of display, for every $j\in[p]$ and $k\in[K]$, we define $\hat{S}_{j}^{(k)}:=\text{supp}\left(\hat{\Theta}_{\setminus j,j}^{(k)}\right)$ as the estimated neighbor set of node $j$ in the $k$-th graph.
@@
Assume Assumptions 1-3 hold and $n\gtrsim\max_{j\in[p]}\left(s_{j}\log p+s_{j}s_{0,j}\log K\right)$. We apply the MIGHT Algorithm 2 with taking
$C_{1}^{\text{alg}}\geq\frac{4\sqrt{3\overline{C}}\left(\kappa/\delta+\sqrt{3}-1\right)}{\kappa/\delta-1},\quad C_{2}^{\text{alg}}\geq\sqrt{\frac{\overline{C}^{2}(\overline{C}^{2}+\underline{C}^{2})}{\underline{C}^{3}}},$
where we define $\delta:=\frac{\overline{C}^{2}}{\overline{C}^{2}+\underline{C}^{2}}\in(0.5,1)$. Then, with probability at least $1-C_{3}e^{-C_{4}\log(pK)}$, we simultaneously have the following properties for every $j\in[p]$:
@@
    $\sum_{k=1}^{K}\frac{n_{k}}{N}\left\|\hat{\Theta}^{(k)}_{\cdot,j}-\Theta^{(k)}_{\cdot,j}\right\|_{2}^{2}\lesssim$ $\frac{K+s_{j}\log p+s_{j}s_{0,j}\log(s_{j}K)}{N}.$ (6)
@@
    Sparse neighbor selection and edge selection: | _k [K] ^S_j^(k) | = O(s_j), and _k[K]| ^S_j^(k) | = O(s_j s_0,j).
The joint estimation bound in (6) consists of three components: (i) estimating the $K$ non-zero diagonal entries $\hat{\Theta}_{j,j}^{(k)}$, with rate $K/N$; (ii) estimating $s_{j}$ neighbors from p1p-1 nodes, with rate $(s_{j}\log p)/N$; and (iii) estimating $s_{j}s_{0,j}$ edges from $s_{j}K$ potential edges, with rate $(s_{j}s_{0,j}\log(s_{j}K))/N$. This decomposition demonstrates that our MIGHT method can adapt to the local $(s_{j},s_{0,j})$ joint structure of each node $j$, yielding an adaptive error bound while ensuring the estimator maintains joint sparsity. Additionally, by combining the node-specific bound (6), we can obtain the joint estimation bound in Frobenius norm as
$\sum_{k=1}^{K}\frac{n_{k}}{N}\left\|\hat{\Theta}^{(k)}-\Theta^{(k)}\right\|_{F}\lesssim\sqrt{\sum_{j\in[p]}\frac{K+s_{j}\log p+s_{j}s_{0,j}\log(s_{j}K)}{N}}.$ (7)
This result offers a more refined error bound than the separation estimation rate $\sqrt{K\cdot\sum_{j\in[p]}\frac{s_{j}\log p}{N}}$. It is also sharper than the bound in existing joint estimation methods Cai et al. (2016); Wang et al. (2017). Furthermore, even when prior support structure information is known, the upper bound derived by Ma and Michailidis (2016), $\sqrt{\frac{\sum_{j\in[p]}s_{j}\log p}{N}}+\sqrt{\frac{K\sum_{j\in[p]}s_{j}}{N}}$, could be less tight than ours in (7) in the case when $\sum_{j\in[p]}s_{j}>\sum_{j\in[p]}s_{j}\frac{s_{0,j}\log(s_{j}K)}{K}$. This scenario can be readily satisfied when the inter-graph local similarity is relatively low for some of the nodes. This demonstrates the superior accuracy and adaptability of our method.